### v2

# Gate-defined single-electron transistors in twisted bilayer graphene

Abstract
Twisted bilayer graphene (tBLG) near the magic angle is a unique platform where the combination of topology and strong correlations gives rise to exotic electronic phases. These phases are gate-tunable and related to the presence of flat electronic bands, isolated by single-particle band gaps. This enables gate-controlled charge confinement, essential for the operation of single-electron transistors (SETs), and allows to explore the interplay of confinement, electron interactions, band renormalisation and the moiré superlattice, potentially revealing key paradigms of strong correlations. Here, we present gate-defined SETs in near-magic-angle tBLG with well-tunable Coulomb blockade resonances. These SETs allow to study magnetic field-induced quantum oscillations in the density of states of the source-drain reservoirs, providing insight into gate-tunable Fermi surfaces of tBLG. Comparison with tight-binding calculations highlights the importance of displacement-field-induced band renormalisation crucial for future advanced gate-tunable quantum devices and circuits in tBLG including e.g. quantum dots and Josephson junction arrays.
twisted bilayer graphene, Coulomb blockade, bias spectroscopy
Figure 1: Device operation and tunable gate-defined conductive channels in tBLG. (a) Scanning force microscopy image of the tBLG device. The relevant gates are highlighted. (b) Schematic cross section of the device along the dashed arrow in panel (a). (c) Tight-binding band structure calculation of tBLG at $\theta=1.25^{\circ}$. The single-particle band gaps $\Delta_{+}$ and $\Delta_{-}$emerging at full filling of the tBLG moiré unit cell are highlighted. (d) Band edge profile along the device showing the formation of the split gate (SG) defined channel by tuning the Fermi level below the SGs into the gap $\Delta_{+}$ and in the channel into the flat band region. (e) Same as in panel (d) but by using the gap $\Delta_{-}$ to define the channel. (f) Two-terminal conductance $G=I/V_{\mathrm{SD}}$ as a function of $V_{\mathrm{BG}}$ and $V_{\mathrm{SG}}=V_{\mathrm{SG1}}=V_{\mathrm{SG2}}$ measured at $V_{\mathrm{SD}}=100\,\mathrm{\mu V}$. The line-cut (right panel) is taken at $V_{\mathrm{SG}}=0.362\,\mathrm{V}$ (white arrow in the conductance map) to compensate for intrinsic doping. Gray arrows mark the half-filling correlated insulating states. (g) Pinch-off traces for gate LG using the gap $\Delta_{-}$ [left panel: at $V_{\mathrm{BG}}=-5\,\mathrm{V}$ and $V_{\mathrm{SG1}}=-2.14\,\mathrm{V}$, $V_{\mathrm{SG2}}=-1.94\,\mathrm{V}$ with $V_{\mathrm{SD}}=100\,\mathrm{\mu V}$; circle in panel (f)] and $\Delta_{+}$ [right panel: at $V_{\mathrm{BG}}=3.5\,\mathrm{V}$ and $V_{\mathrm{SG1}}=V_{\mathrm{SG2}}=3.5\,\mathrm{V}$ with $V_{\mathrm{SD}}=150\,\mathrm{\mu V}$; rhombus in panel (f)]. Insets visualize the band edge bending along the SG-defined channel. (h) Same as in panel (g) but for gate RG.
Single-electron transistors (SETs) [1] are highly sensitive quantum electronic devices that exploit the control of individual charge carriers flowing through a small conducting island or a quantum dot. SETs have been realized in many different material systems (including van der Waals materials such as graphene [2] and 2D semiconductors [3]) and in particular with integrated quantum dots they permit the investigation of band structure properties, the study of electron-electron interactions [4, 5], as well as probing the density of states in nearby reservoirs [6]. All this makes the realisation of tunable SETs in twisted bilayer graphene interesting and promises additional avenues for studying its electronic properties. Twisted bilayer graphene (tBLG) with a twist angle near $1^{\circ}$ not only has a rich phase diagram, including superconductivity [7, 8, 9, 10, 11, 12], strange metal phases and correlated Chern insulators [13, 14, 15, 16, 17], but also offers sizeable single-particle band gaps at finite energy, making charge carrier confinement feasible. The gate-tunability of these phases allowed the demonstration of built-in Josephson junctions [18, 7, 9], superconducting quantum interference devices [19], Aharonov-Bohm rings [20] and superconducting channels [21]. However, well-controlled SETs and quantum dots - all of which require more complex gating structures - have not yet been demonstrated [22].
Here we demonstrate electrostatically controlled charge confinement and SET operation in tBLG based on an advanced gating structure. By using a first set of top gates, we create a narrow conducting channel in the tBLG by exploiting the single-particle band gaps (on the order of several tens of meV) separating the flat bands from the remote conduction and valence band. An additional set of narrow gates across the channel allows local control of the band edge profile and enables the formation of isolated charge islands, very similar as in Bernal-stacked bilayer QD devices [23, 24]. This makes it possible to demonstrate sequential tunneling through flat-band charge islands and using the associated SET as a platform to perform magnetic field spectroscopy measurements, providing information about the Fermi surfaces in the nearby reservoirs and their sensitive dependence on out-of-plane electric fields, confirmed by tight-binding calculations. In addition, our technology provides the basis for future studies of superconducting quantum point contacts, few electron quantum dots, Cooper pair splitters and related quantum devices in the tBLG.
Device operation
The device consists of tBLG with a twist angle close to $\theta\approx 1.25^{\circ}$ (Supporting Section I), which is encapsulated in hexagonal boron nitride (hBN) and placed on a graphite flake acting as a back gate (BG). In addition, there are two metal top gate layers: The first top gate layer consists of a pair of split gates (SGs), SG1 and SG2, which define an $\approx 3\,\mathrm{\mu m}$ long and $\approx 200\,\mathrm{nm}$ wide channel. The second top gate layer (separated from the first gate layer by 20 nm Al2O3) consists of a series of finger gates (FGs) $\approx 200$ nm wide and $\approx 120$ nm apart, placed above the channel so that the FGs cross the channel perpendicular to its orientation (see Methods for fabrication details). In Fig. 1a,b we show a scanning force microscopy image of the device highlighting the left and right FG (LG and RG, respectively), and a schematic cross-section.
To best explain the device operation, we start by discussing the band structure shown in Fig. 1c highlighting the flat bands around zero energy and the single-particle band gaps $\Delta_{+}$ and $\Delta_{-}$ separating them from the remote bands. In a first step, the BG and SG voltages, which locally tune the carrier density $n$, are adjusted so that the Fermi level $E_{\text{F}}$ in the region below the SGs lies in the band gap $\Delta_{+}$ (or $\Delta_{-}$), creating a narrow conducting channel defined by the SGs, as shown in Fig. 1d (or Fig. 1e). In Fig. 1f we show a $V_{\mathrm{BG}}$-$V_{\mathrm{SG}}$ conductance map, where the diagonal features (depending on $V_{\mathrm{BG}}$ and $V_{\mathrm{SG}}$) originate from the areas below the SGs while the horizontal features (independent of $V_{\mathrm{SG}}$) originate from the areas not covered by the SGs. We mark with dashed lines the diagonal features related to the charge neutrality point (CNP, $n=0$ cm-2), the correlated insulating state (CI) at half filling of the fourfold valley and spin degenerate moiré bands, $n=-n_{\mathrm{s}}/2$ [with $n_{\mathrm{s}}=4/$(area of the moiré unit cell)] and the single-particle band gaps $\Delta_{\pm}$ at ±$\pm n_{\mathrm{s}}$ (see labels in Fig. 1f). The diagonal features exhibit a (negative) slope of $\approx 1.5$ in good agreement with the ratio of the tBLG-BG and SG-tBLG capacitances given by $\approx 1.6$ (Supporting Section II and III).
A conductive channel can now be defined by e.g. setting the BG voltage to $V_{\mathrm{BG}}=-5$ V and the SG voltages (slightly asymmetric) to $V_{\mathrm{SG1}}=-2.14$ V and $V_{\mathrm{SG2}}=-1.94$ V (circle in Fig. 1f). This tunes the Fermi level in the region below the SGs in the band gap $\Delta_{-}$, thereby defining a channel as illustrated in Fig. 1e. Next, we can use the individual FGs to completely suppress transport through the channel, as shown in the left panels of Fig. 1g,h. By decreasing one of the FG voltages from $-3.2$ to $-3.6$ V we locally pull the band edges up in energy so that in the region below the LG (Fig. 1g) or the RG (Fig. 1h), the Fermi level is also in the band gap $\Delta_{-}$ (dashed lines in the left insets of Fig. 1g,h).
As the device is gate-tunable, a similar result can be obtained by inverting the gate voltage polarities. This allows for the system to be tuned such that the channel is defined by placing $E_{\text{F}}$ in the gap $\Delta_{+}$ in the region below the SGs (see Fig. 1d); for example, by setting $V_{\mathrm{BG}}=V_{\mathrm{SG}}=3.5\,\mathrm{V}$ (rhombus in Fig. 1f). In this case transport can be completely blocked by using the individual FGs to locally push the band edges down in energy, such that $E_{\mathrm{F}}$ resides in the gap $\Delta_{+}$ (shown by the right panels of Fig. 1g,h).
Figure 2: Formation of a gate-defined SET. (a) Schematic illustration of the band edge profile along the gate-defined channel showing the formation of charge confinement between LG and RG by using the gap $\Delta_{+}$. (b) Large-range charge-stability diagram showing the conductance as a function of LG and RG measured at $V_{\mathrm{BG}}=3.5$ V and $V_{\mathrm{SG}}=3.5$ V (rhombus in Fig. 1d) with $V_{\mathrm{SD}}=100\,\mathrm{\mu V}$. (c) Line-cut showing the onset of Coulomb oscillations along LG for fixed $V_{\mathrm{RG}}=3.12$ V. (d) Zoom-in into the area marked by the black rectangle in panel (b) showing a set of diagonal Coulomb resonances indicating single-electron transport through a charge island located in between the LG and RG. (e) Schematic representation of the SET, with the charge island tunnel-coupled to the source (S) and drain (D) contacts and the corresponding gates highlighted. (f) Finite bias spectroscopy measurements for fixed $V_{\mathrm{RG}}=3.3301$ V [dashed line in panel (d)].
Finally, SET operation can be achieved by using both FGs to modulate the band edge profile along the conductive channel as illustrated in Fig. 2a. Fig. 2b shows a $V_{\mathrm{LG}}$-$V_{\mathrm{RG}}$ conductance map where we observe the transition from high conductance values (regime I, where the FGs are tuned such that we have flat-band carrier transport) to strongly suppressed conductance (regime II, where the FGs push the Fermi level into the gap $\Delta_{+}$). Focusing on a small gate voltage range applied to the LG and RG (black square in Fig. 2b and close-up in Fig. 2d), we observe well-defined conductance oscillations (see also Fig. 2c) that can be tuned equally well by the LG and RG (indicated by the slope of $\approx-1.08$ of the resonances in the $V_{\mathrm{LG}}$-$V_{\mathrm{RG}}$ plane). These periodic conductance oscillations (with a constant spacing of $\Delta V_{\mathrm{LG}}\approx 7.9\,$ mV), which we identify as Coulomb resonances, are a hallmark of the presence of tunneling barriers below both FGs, which consequently electrostatically define a charge island between the left and right FG. The occupation of this charge island, located in the flat bands, can be tuned by the capacitively coupled gates as illustrated in Fig. 2e and confirmed by finite bias spectroscopy measurements revealing characteristic Coulomb diamonds (Fig. 2f).
Highly tunable Coulomb resonances
Figure 3: Fully gate-tunable SET operation in the hole doped regime. (a) Schematic illustration showing the formation of a charge island between LG and RG by using the gap $\Delta_{-}$. (b) Charge-stability diagrams showing diagonal Coulomb resonances for different combinations of BG and SG voltages highlighting the flexibility of the device. Measurements are taken at (i) $V_{\mathrm{BG}}=-3\,\mathrm{V}$ and $V_{\mathrm{SG}}=-3.5\,\mathrm{V}$ at $V_{\mathrm{SD}}=150\,\mu\mathrm{V}$. (ii) $V_{\mathrm{BG}}=-4\,\mathrm{V}$ and $V_{\mathrm{SG}}=-3.23\,\mathrm{V}$ at $V_{\mathrm{SD}}=100\,\mu\mathrm{V}$. (iii) $V_{\mathrm{SG}}=-5\,\mathrm{V}$ and (slightly asymmetric) split gate voltages of $V_{\mathrm{SG1}}=-2.14\,\mathrm{V}$ and $V_{\mathrm{SG2}}=-1.94\,\mathrm{V}$ at $V_{\mathrm{SD}}=100\,\mu\mathrm{V}$ (circle in Fig. 1f) and (iv) $V_{\mathrm{BG}}=-6\,\mathrm{V}$ and $V_{\mathrm{SG}}=-1.7\,\mathrm{V}$ at $V_{\mathrm{SD}}=100\,\mu\mathrm{V}$. (c) Line-cut along the RG extracted from the left most panel in (b) for $V_{\mathrm{BG}}=-3\,\mathrm{V}$ at fixed $V_{\mathrm{LG}}=-4.925$ V showing periodic Coulomb oscillations with an approximate peak spacing of $\Delta V_{\mathrm{RG}}\approx 7.7$ mV. (d) Lower panel: Finite bias spectroscopy measurements as function of $V_{\mathrm{LG}}$ at fixed $V_{\mathrm{BG}}=-5\,\mathrm{V}$ and $V_{\mathrm{RG}}=-3.501$ V (see dashed line in central right panel in (b)). Upper panel: Zero bias Coulomb peak trace extracted from the Coulomb diamond measurement. Note the fact that complete current suppression can be achieved in this regime.
The high tunability of the device is demonstrated by inverting all gate voltages and using the gap $\Delta_{-}$ to form local tunneling barriers essential for the SET operation, as illustrated in Fig. 3a. In Fig. 3b, we show corresponding $V_{\mathrm{LG}}$-$V_{\mathrm{RG}}$ conductance maps for four different $V_{\rm BG}$ values, where $V_{\rm SG}$ is set to ensure that the Fermi level is well within the gap $\Delta_{-}$ in the tBLG region below the SGs, and the LG and RG are fine-tuned to form tunneling barriers below the FGs (see also Supporting Section IV). In each case, we observe periodic Coulomb resonances (diagonal features), demonstrating the versatility of the device to create confinements using both single-particle band gaps of the tBLG electron system to realize a SET. Interestingly, we observe for all investigated regimes a nearly constant spacing of the Coulomb resonances of $\Delta V_{\mathrm{LG}}$ in the range of $(10.6\pm 3.2)$ mV (see e.g. conductance trace shown in Fig. 2c) and $\Delta V_{\mathrm{RG}}$ in the range of $(9.5\pm 2)$ mV, both consistent with the value observed for the electron-doped regime. This, in conjunction with the well-developed diagonal slopes with values in the range of $\approx-0.6$ to $\approx-1.1$ (Supporting Section II) and the possibility to completely suppress the conductance in certain gate voltage regimes, which results in the formation of well-developed Coulomb peaks (Fig. 3d), provides compelling evidence that the island is indeed formed consistently in the channel in between the two FGs. From the characteristic spacing of the Coulomb oscillations $\Delta V_{\mathrm{FG}}$ we can estimate the capacitance between the island and the FGs given by $C_{\mathrm{FG}}=e/\Delta V_{\mathrm{FG}}$ [25] resulting in $C_{\mathrm{FG}}$ in the range of $12$ to $22$ $\mathrm{aF}$. Interestingly, this value is more than a magnitude smaller than the self-capacitance extracted from the charging energy of the islands, which is consistently found to be in the range of $\approx 0.5$ meV (white dashed lines in Fig. 2f and lower panel in Fig. 3d), resulting in a self-capacitance of $C=e^{2}/E_{\mathrm{c}}\approx 320\,\mathrm{aF}$.
Further insights can be gained by comparing this capacitance to the expected BG capacitance, which is expected to provide a dominant contribution. Assuming a simple plate capacitor with the area given by gate spacings ($A=200\times 120$ nm2) we can estimate the BG capacitance by $C_{\mathrm{BG}}=\varepsilon_{0}\varepsilon_{\mathrm{r}}A/d_{\mathrm{b}}$ resulting in $\approx 18$ aF (here we used $d_{\mathrm{b}}\approx 40$ nm and $\varepsilon_{\mathrm{r}}\approx 3.4$). This is also more than an order of magnitude smaller than the self-capacitance $C$ and interestingly comparable to the significantly more distant FG capacitances, strongly indicating that the area of the charge island extends substantially underneath the FGs and (possibly the SGs) resulting in significantly larger island areas, consistent with the large self-capacitance. Moreover, it is expected that the capacitances to the source and drain reservoirs as well as to the SGs will play an important role in this system.
Figure 4: Out-of-plane magnetic field dependence of Coulomb resonances at different BG voltages. (a) Magnetospectroscopy of selected Coulomb resonances for $V_{\mathrm{BG}}=-5\,\mathrm{V}$, $V_{\mathrm{SG1}}=-2.14\,\mathrm{V}$ and $V_{\mathrm{SG2}}=-1.94\,\mathrm{V}$ [see circle in Fig. 1f and compare with Fig. 3b] at $V_{\mathrm{SD}}=220\,\mu\mathrm{V}$ with $V_{\mathrm{RG}}=-3.472$ V. (b) Additional magnetospectroscopy data taken in the other BG regimes shown in Fig. 3b. Measurements are taken at (i) $V_{\mathrm{BG}}=-3\,\mathrm{V}$ and $V_{\mathrm{SG}}=-3.5\,\mathrm{V}$ at $V_{\mathrm{SD}}=220\,\mu\mathrm{V}$ with $V_{\mathrm{LG}}=-4.925$ V. (ii) $V_{\mathrm{BG}}=-4\,\mathrm{V}$ and $V_{\mathrm{SG}}=-3.23\,\mathrm{V}$ at $V_{\mathrm{SD}}=220\,\mu\mathrm{V}$ with $V_{\mathrm{RG}}=-3.79$ V. (iii) $V_{\mathrm{BG}}=-6\,\mathrm{V}$ and $V_{\mathrm{SG}}=-1.7\,\mathrm{V}$ at $V_{\mathrm{SD}}=220\,\mu\mathrm{V}$ with $V_{\mathrm{LG}}=-2.017$ V. For the regimes (ii) and (iii) we performed a background subtraction to highlight the Coulomb resonances. The conductance limits in regime (ii) are scaled by a factor $0.5$ for clarity. (c) Oscillating behavior of $\mathrm{d}V_{\mathrm{LG/RG}}/\mathrm{d}B$ as a function of $1/|B|$ extracted from the tracked Coulomb resonances (i.e. peaks) in panels (a) and (b) averaged over the visible charge transitions. (g) Averaged frequency spectrum of the individual fast Fourier transforms (FFTs) of the oscillations shown in panel (c). For more information see text.
Detection of quantum oscillations
Figure 5: Gate-tunable Fermi surface. (a) $B$-field derivative of the bulk conductance as a function of $V_{\mathrm{BG}}$ (bottom axis) or filling factor $\nu$ (top axis) and out-of-plane magnetic field. The dashed lines highlight the Landau levels emerging from CI at $\nu=-2$ (black) and the CNP at $\nu=0$ (yellow). (b) FFT of the $B$-field depnden quantum oscillations extracted for the different BG voltages (see Fig. 4d). (c) Quantum oscillation frequency $F_{B}$ obtained from tight-binding calculations for displacement field strengths of $D/\varepsilon_{0}=0$ V/nm (orange data) and $D/\varepsilon_{0}=0.48$ V/nm (dark red data). Note, that for converting $\nu$ to $V_{\mathrm{BG}}$ we neglect here stray fields from other surrounding gates. (d) Contour plots of the Fermi surface for different filling factors (i.e. energies) around the $K_{\mathrm{M}}$ points without (top row) and with (bottom row) an applied displacement field. (e) Schematic representation of the $D$-field induced energy shift. For more information see text.
We now turn to the out-of-plane magnetic field ($B$-field) dependence of the Coulomb resonances. In Fig. 4a,b we present Coulomb resonance measurements as a function of $V_{\mathrm{LG}}$ (or $V_{\mathrm{RG}}$) and $B$-field for different BG and SG voltage regimes as shown in Fig. 3b. The dependence of the Coulomb resonances on the $B$-field is generally quite complex. However, there are some common features, including (i) almost no $B$-field dependence for small $B$-fields (up to $0.3-1.2$ T depending on $V_{\mathrm{BG}}$), (ii) a general trend of shifting the Coulomb resonances towards lower energies, i.e. towards the CNP, consistent with previous work on graphene quantum dots [26, 27, 28] and (iii) the appearance of well-defined oscillations at higher $B$-fields, as also recently reported in Bernal-stacked bilayer graphene quantum dots [6] and twist-angle domain-defined charge islands in tBLG [29]. Note that, as in Ref. [29], the spacing between the Coulomb resonances remains constant – within our measurement resolution.
In the following we focus on the observed oscillations as they provide access to interesting information about the electronic structure of the gated flat-band system. Note that these oscillations exhibit a periodicity inversely proportional to the applied $B$-field (see Fig. 4c) and are a consequence of the oscillating Fermi level in the source-drain reservoirs due to the formation of Landau levels [6, 30]. The frequency of the observed oscillations, which have the same origin as Shubnikov-de Haas oscillations, contains valuable information about the Fermi surface $\mathcal{S}$, which can be extracted using the Onsager relation $\mathcal{S}=2\pi eF_{B}/\hbar$, where $F_{B}$ is the frequency of the extracted $1/|B|$ oscillations. To study the Fermi surface and its dependence on band filling, we conducted measurements of the $B$-field dependence of the Fermi surface for different values of $V_{\mathrm{BG}}$, as demonstrated in Fig. 4a,b. In Fig. 4c, we present the corresponding $B$-field derivatives of the gate voltage peak positions, plotted as a function of inverse $B$-field. Each trace represents an average taken over at least three Coulomb resonances (Supporting Section V). The corresponding fast Fourier transforms (FFTs) are shown in Fig. 4d, which highlight that there is indeed one dominant $1/|B|$ frequency, $F_{B}$. This frequency shifts to lower values when $V_{\mathrm{BG}}$ is decreased.
Interestingly, there are some at first sight unexpected differences when comparing this data with the bulk magneto-conductance measurements for which the SGs and FGs are tuned such that the entire tBLG area is uniformly conductive (Supporting Section I). In Fig. 5a we show such a bulk measurement, specifically $\mathrm{d}G/\mathrm{d}B$ as function of $V_{\mathrm{BG}}$ and $B$-field and find the well expected emerging Landau fan from $V_{\mathrm{BG}}=0$ V (see yellow dashed lines) which is connected to a Fermi surface area $\mathcal{S}$ that vanishes at the CNP (filling factor $\nu=0$) in good agreement with earlier work [18]. Moreover, we observe additional fans emerging from the CI states at $\nu=\pm 2$ (see e.g. black dashed lines highlighting the fan at $\nu=-2$) implying that $\mathcal{S}$ should exhibit a minimum (if not vanishing) at these points. Let us now turn to the BG voltage dependent Fermi surface, i.e. $F_{B}$, extracted from the Coulomb resonance measurements discussed in Fig. 4 and summarised in Fig. 5b. Here we observe the expected linear tuning of the main $1/|B|$ frequency component, but there is no minimum (or zero) at $V_{\mathrm{BG}}\approx-4$ V, where we would expect the CI state at $\nu=-2$ (compare to Fig. 5a). Furthermore, by extending a linear slope to $V_{\mathrm{BG}}=0\,\mathrm{V}$ (dashed line in Fig. 5b), we conclude that there is most likely a finite Fermi surface at $V_{\mathrm{BG}}=0$ V, in contrast to the measurements shown in Fig. 5a, and in contrast to large-scale tight-binding calculations (see orange trace in Fig. 5c) [31, 32, 33, 34, 35]. Note that our tight-binding calculations (details in Supporting Section VII) are in good agreement with previous works [18, 36] and our bulk measurements (Fig. 5a). However, when compared with the data obtained from Coulomb resonance magnetospectroscopy (Fig. 5b), the tight-binding calculation reveals an additional difference: the expected Lifshitz transition at $\nu\approx-1.5$ (black arrow in Fig. 5c – expected at $V_{\mathrm{BG}}\approx-3$ V), which leads to a sign reversal of the slope of $F_{\mathrm{B}}$ vs. $V_{\mathrm{BG}}$ is completely absent in the data shown in Fig. 5b. This Lifshitz transition, which is related to a van Hove singularity in the density of states, describes the transition from Fermi surfaces associated with valley-exclusive hole-like orbits around the moiré Dirac points $K_{\mathrm{M}}^{(\prime)}$ to Fermi surfaces associated with electron-like orbits centred around $\Gamma_{\mathrm{M}}$, as shown in Fig. 5d (upper row).
The absence of the Lifshitz transition and the indication of a finite Fermi surface at the expected CNP can in principal be explained by the presence of (i) an out-of-plane electric field [37] or (ii) heterostrain [38, 39], both having sizeable impact on the band structure of tBLG near the magic angle. From the bulk measurements (Fig. 5a), in particular from the fact that there is only one Landau fan emerging from the CNP, we rule out heterostrain as the dominant cause of the finite electron/hole Fermi surfaces at the CNP [40], making out-of-plane displacement fields ($D$-fields) the prime candidate responsible for the band deformations. Such $D$-fields are likely present due to the multiple gate layers with different potentials, i.e. gate voltages. Crucially, a finite $D$-field leads to an energy difference between the $p_{\mathrm{z}}$-orbitals in the upper and lower graphene layer of the tBLG, which creates an energetic asymmetry between the cones (band touching points) at the moiré Dirac $K_{\mathrm{M}}$ and $K_{\mathrm{M}}^{\prime}$ points. This energetic offset induced by the displacement field (Fig. 5e) results (i) in a finite Fermi surface at the CNP, where both conductance and valence band carriers are contributing, and (ii) in a shift of the Lifshitz transition to higher hole filling.
In Fig. 5c (dark red data points) and Fig. 5d (lower row), we present the corresponding tight-binding calculations for an applied displacement field of $D/\varepsilon_{0}=0.48$ V/nm, which is a reasonable value for the applied gate voltages (Supporting Section IX). The energetic offset of the Dirac cones at the $K_{\mathrm{M}}$ and $K_{\mathrm{M}}^{\prime}$ points results in a finite Fermi surface (i.e. in a finite $F_{B}$) at the CNP in quantitative agreement with the experiment (compare Fig. 5b,c). Near the CNP, the Fermi surface within a single valley is made up of valence band (orange) and conductance band (blue) pockets as shown in the lower left panel in Fig. 5d. With increasing hole filling the imbalance between the valance band and conduction band pockets increases (Fig. 5d) with the overall consequence that the Lifshitz transition is shifted to values below $\nu<-3$ in agreement with the experimental data (compare Fig. 5b,c). Moreover, we expect that the presence of strong displacement fields in the gate-defined device geometry significantly screens the long-ranged Coulomb interaction [41, 11, 42] suppressing the formation of the correlated insulating states at $\nu=-2$, in agreement with our experimental data (Fig. 5b) and in line with theoretical predictions of interaction-induced band renormalisation [43, 44, 45, 46] (Supporting Section VIII). Thus, our findings stress the significant impact of out-of-plane electrical fields on the band structure of near magic-angle tBLG.
Discussion and conclusions
In conclusion, our measurements show that we can implement gate-defined SET in tBLG and use them as a platform to gain insights in band deformations of (near) magic angle tBLG in the presence of out-of-plane electric fields. Essentially, our extracted data fill the gap between the large-angle regime ($\theta\gg 1^{\circ}$), which is characterized by large bandwidths and comparatively little sensitivity to external electrical fields [40, 18] and the small-angle regime ($\theta\ll 1^{\circ}$), where the presence of electric fields causes the formation of topological helical networks [47, 48, 49]. An important observation is the absence of the correlated insulating state in Fig. 5b, while it is prominently present in the bulk transport data shown in Fig. 5a.
Given the recent theoretical and experimental study on the importance of long-ranged Coulomb interactions in the formation of correlated states in (near) magic-angle tBLG [50, 11, 42], we emphasize the flexibility of our device architecture to study electric field-induced screening of electronic interactions. In this respect, future experiments even closer to the magic angle could provide much-needed answers to a number of pressing questions related to correlated phenomena such as superconductivity and strange metal phases and their interplay with topology. In addition, the technology presented can enable the development of complex quantum devices and circuits based on tBLG, ranging from coupled quantum dots, Cooper-pair splitters, chains of Josephson junction-based SQUIDs to single-photon detector arrays.
Acknowledgements This project has received funding from the European Research Council (ERC) under grant agreement No. 820254, the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through SPP 2244 (Project No. 535377524) and under Germany’s Excellence Strategy - Cluster of Excellence Matter and Light for Quantum Computing (ML4Q) EXC 2004/1 - 390534769, the FLAG-ERA grants No. 437214324 TATTOOS, No. 471733165 PhotoTBG, No. 534269806 ThinQ by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), and by the Helmholtz Nano Facility [51]. K.W. and T.T. acknowledge support from the JSPS KAKENHI (Grant Numbers 21H05233 and 23H02052) and World Premier International Research Center Initiative (WPI), MEXT, Japan.
Author contributions C.S. and A.R. conceived the experiment. A.R. built the device and performed the measurements. A.F. performed the theoretical simulations. A.R. and A.A. performed the data analysis under the supervision of C.S.. E.I., K.H and L.B. provided assistance with the measurement setup. M.O. performed the ALD step. S.T. and F.L. performed the electron beam lithography. K.W. and T.T. supplied the hBN crystals. A.R., A.F., R.J.D., D.M.K., B.B. and C.S. discussed the experimental and theoretical data and wrote the manuscript.
Data availability The data supporting the findings of this study are available in a Zenodo repository under [insert DOI link here].
Methods
Device Fabrication
The device consists of a hBN-tBLG-hBN heterostructure (thicknesses of the top and bottom hBN are measured via atomic force microscopy and approximately given by $d_{\mathrm{t}}\approx 25$ nm and $d_{\mathrm{b}}\approx 40$ nm, respectively) with a graphite BG fabricated by a ’laser-cut-and-stack’ technique following a standard PDMS/PC transfer process [52, 53, 54]. After one-dimensional side contacting (Cr/Au, $5/50$ nm) of the tBLG [55], we fabricate the first gating layer containing the SGs (Cr/Au, $5/50$ nm) on top of the upper hBN layer. These gates are $3$ $\mathrm{\mu}$m wide and are separated by $200$ nm, defining a quasi one-dimensional channel [see scanning force microscopy image in Fig. 1a]. Following the SG metallisation, we evaporate a hard mask (Al, $60$ nm) to subsequently structure the mesa by means of reactive ion etching ($\mathrm{SF_{6}/O_{2}}$). In the next step, we use atomic layer deposition (ALD) of alumina ($\mathrm{Al_{2}O_{3}}$) to make a $15$ nm thick insulating layer separating the SGs from the FG layer. Finally, we structure the FGs (Cr/Au, $5/170$ nm), which intersect the previously defined channel. The three FGs are approximately $200$ nm wide and are separated by around $120$ nm. A schematic cross section along the device is shown in Fig. 1b
Measurement Setup
All measurements presented in this manuscript were performed in a ${}^{3}\mathrm{He}/{}^{4}\mathrm{He}$ (Oxford Instruments, Triton 200) dilution refrigerator at a temperature below $100$ mK. In a two-terminal setup we apply a fixed bias voltage over a home-built IV-converter via the source-drain contacts highlighted in Fig. 1a and parallel measure the current with a digital multimeter (Agilent 34401). For the bias spectroscopy measurements shown in Fig. 2f and Fig. 3d we used a lock-in amplifier (Stanford Research SR830) at a frequency of $14.5$ Hz in a two-terminal setup.